# Federated Multi-Agent Mapping for Planetary Exploration

### v3

@@ I INTRODUCTION
In space exploration, the traditional paradigm of sending a single rover or spacecraft is giving way to new architectures that allow for faster, more efficient, and autonomous solutions that enable greater scientific returns and understanding. For example, the Martian Ingenuity helicopter scouted terrain in advance for the Perseverance rover with the aim of identifying scientific targets of interest in advance and allowing for faster traverses for the rover [1]. Indeed, such multi-agent mission concepts where teams of rovers or drones collaborate are poised to play a greater role in future exploration to more remote and far-flung planetary bodies in the Solar System [2, 3]. A pressing challenge in enabling such mission architectures is that of information sharing, where the traditional concept-of-operations involves sending data generated on-board back to ground station operators on Earth to act upon. This approach scales poorly with an increasing number of agents and can lead to costly time delays due to limited data transfer rates, both problems exacerbated by the increasing use of autonomy in planetary exploration [4].
In this work, we draw inspiration from the field of  federated learning (FL) and demonstrate how FL can enable teams of planetary robots to effectively carry out science and exploration tasks in the presence of such operating constraints. FL is a learning-based approach in which statistical models are trained in a distributed fashion without transmitting data directly [5, 6]. Although FL has become popular in medical and consumer applications due privacy concerns that prevent the direct sharing of sensitive data, it has received limited attention for applications in space robotics and planetary exploration. As such, we note that the upcoming  Cooperative Autonomous Distributed Robotic Exploration (CADRE) mission ideally lends itself to the infusion of FL for the task of multi-agent distributed mapping. CADRE is a Lunar rover mission in which a team of three Lunar rovers (Figure 2) map the Lunar environment by using distributed ground penetrating radar instruments [7]. During the Exploration phase of the mission, the three rovers will compute traversability maps on-board and transmit the raw map data to a central base station, where a global map of the environment is reconciled for use in science planning.
@@
Figure 2: CADRE rovers in a Clean Room [13].
@@ II-A Autonomy in Planetary Exploration
However, in a cooperative multi-agent mission, each individual agent’s map offers limited situational awareness for use in science and mission planning. Instead, map information must be shared between the agents to construct a global map, whereby valuable scientific measurement targets can be identified. For CADRE, this global map construction is accomplished through the use of the MoonDB framework [16], which is a centralized database on the mission base station that stores and fuses the traversability maps generated by each agent. The key pitfall of MoonDB is that as the number of agents grow or each agent explores, the volume of data generated grows intractably and this bottlenecks global map construction.
@@ II-B Neural and Distributed Mapping
NERFs [8] offer a way to create 3D representations from 2D images using deep neural networks to model volumetric density and color. A NeRF function $F_{\theta}:(x,d)\rightarrow(c,\sigma)$ takes a 3D location $(x,y,z)$ and viewing direction $d$ as input, outputting radiance (color $c$) and density $\sigma$. While NeRFs are used for 3D scene representation, a 2D version, also known as coordinate-based MLPs [19], applies similar principles to map image coordinates $(x,y)$ to color value $c$ (RGB).
In neural mapping, networks are biased toward learning low-frequency features, limiting their ability to capture high-frequency details like sharp textures [20]. To address this, mapping inputs to higher-dimensional spaces improves the modeling of high-frequency variations [21, 8]. Approaches like RNR-Map [22] use NeRFs for 3D environment embedding, aiding localization and navigation, but struggle with unseen or ambiguous areas.
Federated learning approaches presented in [23, 24] reduce communication and training costs by collaboratively training NeRFs across distributed devices. However, challenges remain in handling non-IID data, integrating diverse sensors, and adapting to dynamic environments, which are critical for multi-agent planetary exploration.
@@ III-A Overview
To address the challenges in Section  II, we introduce a federated learning framework for multi-agent planetary exploration. Each agent learns local maps using 2D networks, with offline meta-training on Earth-based data. This establishes a robust prior that enables networks to rapidly learn from limited local data and viewpoints, addressing the challenges of non-IID data and ensuring local model performance, while supporting heterogeneous sensor configurations. Summarizing our contributions:
@@
Our map refinement process uses morphological image processing (e.g. connected elements) to fill gaps and remove noise artifacts, improving map clarity. Fig. 3 illustrates the approach, and Alg. 1 provides the pseudocode.
@@ III-D Network Definition
The network is defined as a simple convolutional architecture for 2D implicit mapping, with the default configuration consisting of four convolutional layers (kernel size 1, no padding) and an output of 3 (RGB), using ReLU activations [28] and batch normalization. This configuration functions like a fully-connected (dense) layer, making the network resemble a multi-layer perceptron (MLP). The architecture is scalable, allowing flexibility from simple linear layers to more complex convolutional ones. The parameter count is fixed, depending only on the input and output channels, not input resolution, making it independent of spatial dimensions. The network uses 256 channels to match positional encoding and is trained using the ADAM optimizer [29] with Mean Squared Error (MSE) loss.
@@ III-E Federated Mapping
Using the network from III-D, our federated mapping process starts with learning an unexplored grid (Line 4) and meta-training (Line 5), to enable swift adaptation. Each agent is initialized with meta-trained parameters (Line 7).
Agents explore independently, generating local maps (Line 9) and training the network using the ADAM optimizer and MSE loss (Line 10). Trained parameters are sent to the central server (Line 11) for aggregation (e.g., FedAvg) to create an updated global model (Line 13), which is redistributed to agents (Line 14). The global model is then used to generate a global map at each agent (Line 15), followed by gap filling and artifact removal (Line 16). Importantly, we focus on few-shot federated learning, aiming for effective map refinement with minimal communication rounds, iterating as agents continue exploring.
@@ III-F Federated Map Refinement
1:$M_{0}$ an empty map, $M_{\textrm{global}}$ the federated global map, $D_{\textrm{trav}}$ traversability map dataset, $F$ global reference frame
@@
4:$\theta_{0}\leftarrow\textrm{Train}(\textrm{Net},M_{0})$
5:$\theta\leftarrow\textrm{MetaTrain}(\textrm{Net},\theta_{0},D_{\textrm{trav}})$
@@
10:     $\theta_{i}\leftarrow\arg\min_{\theta}L(M_{i},\textrm{Net}(\theta))$triangleright Local network training
@@
13:$\theta_{\textrm{global}}\leftarrow\frac{1}{N}\sum_{i=1}^{N}\frac{n_{k}}{n}\theta_{i}$triangleright Federated Aggregation
@@
15:$M_{\textrm{global}}\leftarrow\textrm{Net}(\theta_{\textrm{global}})$triangleright Generate global map
16:$M_{\textrm{global}}\leftarrow\textrm{RefineMap}(M_{\textrm{global}})$triangleright Filter and remove noise
@@ IV-A Datasets
TABLE I: Comparative analysis of initialization methods for map learning. We detail the PSNR after two iterations of test optimization, showing that the meta-learned initialization (Meta) outperforms other methods. Additionally, we present the avg. number of iterations required for the other methods to reach the Meta’s performance level after two iterations.
@@ IV-B Metrics
Additionally, we evaluate path planning by comparing A* paths on true versus learned maps. We assess path viability using precision, recall, and F1 scores, which provide a balanced measure of the model’s utility for autonomous navigation, as seen in prior work [30].
@@ IV-C Meta-initialization Ablation Results
In Fig. 4, we compare meta-initialization with random initialization after 1000 iterations and a learning rate of $1e^{-4}$ across $200$ test samples per run. The meta-initialized network quickly captures fine-grained map details, while the randomly initialized one takes significantly longer to reach similar fidelity. As shown in Table I, the meta-learned network, initiated from an empty map, achieves an average PSNR of 13.30 after just one learning step (two forward passes) on novel test maps, outperforming random initialization (8.16 PSNR) and empty map initialization (9.37 PSNR). Random initialization requires 68.68 ± 1.06 forward passes, and empty map initialization needs 26.98 ± 1.84 to match the meta-trained network’s PSNR, demonstrating faster adaptation. Thus, we see that meta-initialization using readily available Earth-based datasets enables performant FL and fast adaptation for space-based applications.
@@ IV-D Federated Mapping Results
To test this hypothesis, we used two out-of-distribution datasets: the Athabasca Glacier and a simulated Martian surface. As illustrated in Fig. 6, four agents explored distinct areas, learning the generated local maps ($LR=1e^{-4}$), and sharing trained parameters in a single communication round ($R=1$). Table II provides a comprehensive evaluation of one-shot mapping performance across key metrics: PSNR (map reconstruction quality), SSIM (structural similarity) and F1 scores (path planning efficacy), using consistent settings (4 agents, 100 local epochs).
@@
(1) FedAdam integrates Adam’s adaptive learning rates into federated learning but performs poorly in our experiments, with low PSNR, SSIM, and near-zero F1 scores. The main issue is likely its sensitivity to gradient differences across agents. In highly heterogeneous surfaces such as the glacier crevasses, where local data distributions differ significantly, the adaptive adjustments seem to have caused instability during aggregation, leading to excessive or insufficient learning rate adjustments for effective global models. (2) FedYogi, an adaptation of the Yogi optimizer for federated settings, improves over FedAdam, especially on the Mars dataset with an F1 score of 0.87. However, it still struggled with the high data heterogeneity, as the learning rate adjustments were still too reactive to the large variations in local environments. (3) FedAvg outperforms both adaptive methods, achieving the highest PSNR (10.85), SSIM (0.47), and F1 scores (0.95 on Athabasca Glacier, 0.94 on Mars). These results suggest that FedAvg’s simpler approach of averaging local updates, without dynamic learning rate adjustments, proved more stable in the tested environments, leading to superior map reconstruction and path planning across diverse terrains. In contrast, adaptive methods like FedAdam and FedYogi struggled with instability during model aggregation, making them less effective for federated learning in multi-agent planetary exploration.
(a) Federated mapping on the Athabasca Glacier (black = passable). See Fig. 6 for detailed progression.
(b) Federated mapping on a Martian surface (brown = passable). See Fig. 6 for detailed progression.
Figure 6: Federated map learning results for both the Athabasca Glacier (top row) and a Martian surface (bottom row). For both, the figure progresses as follows: the first four subfigures display the local maps learned by individual agents. Subfigure 5 shows the ground truth global map. Subfigure 6 presents the final global map obtained by combining the agent maps through federated learning.
@@ IV-E Data Transfer Comparison
Transmitting the full model parameters requires approx. 399 KB. This is already more bandwidth-efficient than sending raw map data for any instance of the grayscale, optimal mixture of gaussians (OMG) map proposed for Mars rotorcrafts [34], or the converging covariance map (CCM) used for CADRE [7]. In particular, we see that the CADRE CCM requires approximately 2.88 MB for a 400$\times$400 cell map (18 bytes/cell). We note that this comparison is for a single CCM map, whereas the Exploration phase of the CADRE mission entails computing multiple CCMs as the rover drives and each one of these CCM instances is transmitted back to the base station for map merging. In addition, JPEG with quality factor (Q=26) can perform lossy compression for a raw grayscale map (2000×2000) to 402 kB -nearly matching the full model parameters. In this configuration, it can yield high quality metrics (PSNR: 25.88, SSIM: 0.90), and a similar downstream path planning F1 score of 0.96.
However, our approach stands apart from conventional image compression like JPEG because the model can leverage gradient sparsification. Techniques such as top-k selection, as demonstrated in Deep Gradient Compression (DGC) [35], can reduce the data transmitted by roughly 99% compared to full gradient updates for CNN-based ResNet architectures. In our task, where the full model is approximately 399 kB, this implies that only about 1%—roughly 4 kB—needs to be sent per update round. This dramatic reduction in update size significantly outperforms traditional JPEG compression. Moreover, the meta-initialization strategy functions similarly to the warm-up training phase in DGC. This warm-up establishes a strong prior, such that that early updates are less noisy and more stable. Given the relatively stable nature of space environments, where maps exhibit only gradual changes, our network could produce extremely sparse incremental updates.
@@ V Conclusion
In this paper, we introduce a federated multi-agent mapping strategy for space exploration robotics, utilizing implicit neural mapping for efficient and detailed map creation. This method supports rapid adaptation in diverse environments through meta-initialization and collaborative learning. Validation on mission-like datasets confirms its effectiveness. In future work, we aim to apply FL for mission concepts in which information must be shared across heterogeneous agents (e.g., a rover and a helicopter). Finally, we plan to extend our approach to 3D mapping that capture richer scene representations beyond traversability maps, further enhancing the compression benefits.
@@ ACKNOWLEDGMENT
The research was carried out in part at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with NASA (80NM0018D0004). The authors thank Jean-Pierre de la Croix, Georgios Georgakis, Shehryar Khattak, Mike Paton, Jakob Eg Larsen and Niels H. Pontoppidan for their discussions during the development of this work.

### v4

@@ I INTRODUCTION
In space exploration, the traditional paradigm of sending a single rover or spacecraft is giving way to new architectures that allow for faster, more efficient, and autonomous solutions that enable greater scientific returns and understanding. For example, the Martian Ingenuity helicopter scouted terrain in advance for the Perseverance rover with the aim of identifying scientific targets of interest in advance and allowing for faster traverses for the rover [2]. Indeed, such multi-agent mission concepts where teams of rovers or drones collaborate are poised to play a greater role in future exploration to more remote and far-flung planetary bodies in the Solar System [21]. A pressing challenge in enabling such mission architectures is that of information sharing, where the traditional concept-of-operations involves sending data generated on-board back to ground station operators on Earth to act upon. This approach scales poorly with an increasing number of agents and can lead to costly time delays due to limited data transfer rates, both problems exacerbated by the increasing use of autonomy in planetary exploration [27].
In this work, we draw inspiration from the field of  federated learning (FL) and demonstrate how FL can enable teams of planetary robots to effectively carry out science and exploration tasks in the presence of such operating constraints. FL is a learning-based approach in which statistical models are trained in a distributed fashion without transmitting data directly [10, 8]. Although FL has become popular in medical and consumer applications due to privacy concerns that prevent the direct sharing of sensitive data, it has received limited attention for applications in space robotics and planetary exploration. As such, we note that the upcoming  Cooperative Autonomous Distributed Robotic Exploration (CADRE) mission ideally lends itself to the infusion of FL for the task of multi-agent distributed mapping. CADRE is a Lunar rover mission in which a team of three Lunar rovers (Figure 2) map the Lunar environment by using distributed ground penetrating radar instruments [3]. During the Exploration phase of the mission, the three rovers will compute traversability maps on-board and transmit the raw map data to a central base station, where a global map of the environment is reconciled for use in science planning.
@@
Figure 2: The CADRE mission, with the flight models shown here being tested in a clean room, is an upcoming multi-agent Lunar rover mission featuring three rovers performing distributed measurements of the Lunar surface and sub-surface, that aims to demonstrate how autonomy can dramatically transform scientific exploration capabilities across the Solar System [3].
@@ II-A Autonomy in Planetary Exploration
However, in a cooperative multi-agent mission, the situational awareness provided by each individual agent’s map is limited for science and mission planning. Instead, map information must be shared between the agents to construct a global map, whereby valuable scientific measurement targets can be identified. For CADRE, ground station operators specify a region to be explored, after which exploration is performed autonomously. Next, frontier candidates are computed from the shared traversability map and rovers are dynamically assigned to targets that minimize cost and maintain communication with the base. This global map construction is accomplished using MoonDB [20], a centralized database on the mission base station that stores and fuses the traversability maps generated by each agent. The key pitfall of MoonDB is that as the number of agents grow and each agent explores, the volume of data generated grows intractably and this bottlenecks global map construction.
@@ II-B Neural and Distributed Mapping
NERFs [12] offer a way to create 3D representations from 2D images using deep neural networks to model volumetric density and color. A NeRF function $F_{\theta}:(x,d)\rightarrow(c,\sigma)$ takes a 3D location $(x,y,z)$ and viewing direction $d$ as input, outputting radiance (color $c$) and density $\sigma$. While NeRFs are used for 3D scene representation, a 2D version known as coordinate-based MLPs [22] applies similar principles to map image coordinates $(x,y)$ to color value $c$ (RGB).
In neural mapping, networks are biased toward learning low-frequency features, limiting their ability to capture high-frequency details such as sharp textures [18]. To address this, mapping inputs to higher-dimensional spaces improves the modeling of high-frequency variations [25, 12]. Approaches such as RNR-Map [7] use NeRFs for 3D environment embedding, aiding localization and navigation, but struggle with unseen or ambiguous areas. The federated learning approaches presented in [5, 23] reduce communication and training costs by collaboratively training NeRFs across distributed devices. However, challenges remain in handling non-IID data, integrating diverse sensors, and adapting to dynamic environments, all of which are critical for multi-agent planetary exploration.
@@ III-A Overview
To address the challenges described in Section  II, we introduce a federated learning framework for multi-agent planetary exploration. Each agent learns local maps using 2D networks, with offline meta-training on Earth-based data. This establishes a robust prior that enables networks to rapidly learn from limited local data and viewpoints, addressing the challenges of non-IID data and ensuring local model performance, while supporting heterogeneous sensor configurations. Summarizing our contributions:
@@
Our map refinement process uses morphological image processing (e.g. connected elements) to fill gaps and remove noise artifacts, improving map clarity. Fig. 3 illustrates the approach and Alg. 1 provides high-level pseudocode.
@@ III-D Network Definition
The network is defined as a simple convolutional architecture for 2D implicit mapping, with the default configuration consisting of four convolutional layers (kernel size 1, no padding) and an output of 3 (RGB), using ReLU activations [1] and batch normalization. This configuration functions similar to a fully-connected (dense) layer, making the network resemble a multi-layer perceptron (MLP). The architecture is scalable, allowing flexibility from simple linear layers to more complex convolutional ones. The parameter count is fixed, depending only on the input and output channels, not input resolution, making it independent of spatial dimensions. The network uses 256 channels to match positional encoding and is trained using the ADAM optimizer [6] with Mean Squared Error (MSE) loss.
@@ III-E Federated Mapping
Using the network from III-D, our federated mapping process starts with learning an unexplored grid (Line 4) and meta-training (Line 5), to enable swift adaptation. Each agent is initialized with meta-trained parameters (Line 7). Agents explore independently, generating local maps (Line 9) and training the network using the ADAM optimizer and MSE loss (Line 10). Trained parameters are sent to the central server (Line 11) for aggregation (e.g., FedAvg) to create an updated global model (Line 13), which is redistributed to agents (Line 14). The global model is then used to generate a global map at each agent (Line 15), followed by gap filling and artifact removal (Line 16). Importantly, we focus on few-shot federated learning, aiming for effective map refinement with minimal communication rounds, iterating as agents explore.
@@ III-F Federated Map Refinement
[…]
@@
[…]
@@
10:  $\theta_{i}\leftarrow\arg\min_{\theta}L(M_{i},\textrm{Net}(\theta))$ $\triangleright$ Local network training
@@
[…]
@@
[…]
@@ IV-A Datasets
TABLE I: Comparative analysis of initialization methods for map learning. We report the PSNR after two iterations of test optimization, showing that the meta-learned initialization (Meta) outperforms other methods. Additionally, we present the avg. number of iterations required for the other methods to reach the Meta’s performance level after two iterations.
@@ IV-B Metrics
We evaluate path planning by comparing A* paths on true versus learned maps. A true positive is counted when the paths are identical. A false positive occurs when the learned map yields a path where none exists in the ground truth and a false negative when it fails to find one that does. This strict setup tests whether the map supports the same optimal path, offering a strong measure of navigational fidelity, similar to prior work [17].
@@ IV-C Meta-initialization Ablation Results
In Fig. 4, we compare meta-initialization with random initialization after 1000 iterations and a learning rate of $10^{-4}$ across $200$ test samples per run. The meta-initialized network quickly captures fine-grained map details, while the randomly initialized one takes significantly longer to reach similar fidelity. As shown in Table I, the meta-learned network, initiated from an empty map, achieves an average PSNR of 13.30 after just one learning step (two forward passes) on novel test maps, outperforming random initialization (8.16 PSNR) and empty map initialization (9.37 PSNR). Random initialization requires 68.68 ± 1.06 forward passes, and empty map initialization needs 26.98 ± 1.84 to match the meta-trained network’s PSNR, demonstrating faster adaptation. Thus, we see that meta-initialization using readily available Earth-based datasets enables performant FL and fast adaptation for space-based applications.
@@ IV-D Federated Mapping Results
(a) Federated mapping on the Athabasca Glacier. Progression described in the figure caption.
(b) Federated mapping on a Martian surface.
Figure 6: Federated map learning results for both the Athabasca Glacier (top row) and a Martian surface (bottom row). For both, the figure progresses as follows: the first four subfigures display the local maps learned by individual agents. Subfigure 5 shows the ground truth global map. Subfigure 6 presents the final global map obtained by combining the agent maps through federated learning. Green = free; Red = impassable; Gray = unknown.
To test this hypothesis, we used two out-of-distribution datasets: the Athabasca Glacier and a simulated Martian surface. As illustrated in Fig. 6, four agents explored distinct areas, learning the generated local maps and sharing trained parameters in a single communication round ($R=1$). Table II provides a comprehensive evaluation of one-shot mapping performance across key metrics: PSNR (map reconstruction quality), SSIM (structural similarity) and F1 scores (path planning efficacy), using consistent settings (4 agents, 100 local epochs).
@@
(1) FedAdam integrates Adam’s adaptive learning rates into federated learning but performs poorly in our experiments, with low PSNR, SSIM, and near-zero F1 scores. The main issue is likely its sensitivity to gradient differences across agents. In highly heterogeneous surfaces such as the glacier crevasses, where local data distributions differ significantly, the adaptive adjustments seem to have caused instability during aggregation, leading to excessive or insufficient learning rate adjustments for effective global models. (2) FedYogi, an adaptation of the Yogi optimizer for federated settings, improves over FedAdam, especially on the Mars dataset with an F1 score of 0.87. However, it still struggled with the high data heterogeneity, as the learning rate adjustments were still too reactive to the large variations in local environments. (3) FedAvg outperforms both adaptive methods, achieving the highest PSNR (10.85), SSIM (0.47), and F1 scores (0.95 on Athabasca Glacier, 0.94 on Mars). These results suggest that FedAvg’s simpler approach of averaging local updates, without dynamic learning rate adjustments, proved more stable in the tested environments, leading to superior map reconstruction and path planning across diverse terrains. In contrast, adaptive methods such as FedAdam and FedYogi struggled with instability during model aggregation, rendering them less effective for federated learning in multi-agent planetary exploration.
@@ IV-E Data Transfer Comparison
Transmitting the full model parameters requires approx. 399 KB. This is already more bandwidth-efficient than sending raw map data for any instance of the grayscale, optimal mixture of gaussians (OMG) map proposed for Mars rotorcrafts [16], or the converging covariance map (CCM) used for CADRE [3]. In particular, we see that the CADRE CCM requires approximately 2.88 MB for a $\times400$ cell map (18 bytes/cell). We note that this comparison is for a single CCM map, whereas the Exploration phase of the CADRE mission entails computing multiple CCMs as the rover drives and each one of these CCM instances is transmitted back to the base station for map merging. In addition, JPEG with quality factor (Q=26) can perform lossy compression for a raw grayscale map (2000×2000) to 402 kB -nearly matching the full model parameters. In this configuration, it can yield high quality metrics (PSNR: 25.88, SSIM: 0.90), and a similar downstream path planning F1 score of 0.96.
However, our approach stands apart from conventional image compression such as JPEG because the model can leverage gradient sparsification. Techniques such as top-k selection, as demonstrated in Deep Gradient Compression (DGC) [9], can reduce the data transmitted by roughly 99% compared to full gradient updates for CNN-based ResNet architectures. In our task, where the full model is approximately 399 kB, this implies that only about 1%—roughly 4 kB—needs to be sent per update round. This dramatic reduction in update size significantly outperforms traditional JPEG compression. Moreover, the meta-initialization strategy functions similarly to the warm-up training phase in DGC. This warm-up establishes a strong prior, such that early updates are less noisy and more stable.
@@ V Conclusion
In this paper, we introduce a federated multi-agent mapping strategy for space exploration robotics, utilizing implicit neural mapping for efficient and detailed map creation. This method supports rapid adaptation in diverse environments through meta-initialization and collaborative learning. Validation on mission-like datasets confirms its effectiveness. In future work, we aim to apply FL for mission concepts in which information must be shared across heterogeneous agents (e.g., a rover and a helicopter). Finally, we plan to extend our approach to 3D mapping that capture richer scene representations beyond traversability maps.
@@ ACKNOWLEDGMENT
The authors thank Preston Culbertson, Jean-Pierre de la Croix, Georgios Georgakis, Shehryar Khattak, Jakob Eg Larsen, Mike Paton, and Niels H. Pontoppidan for their discussions during the development of this work.